Wheat is produced under perfectly competitive conditions. Individual wheat farmers have U-shaped, long-run average cost curves that reach a minimum average cost of $3 per bushel when 1,000 bushels are produced.
a) If the market demand curve for wheat is given by QD = 2,600,000 - 200,000P, where Qp is the number of bushels demanded per year and P is the price per bushel, in long-run equilibrium what will be the price of wheat, how much total wheat will be demanded, and how many wheat farms will there be?
1
Expert's answer
2019-07-24T15:53:39-0400

QD = 2,600,000 - 200,000P

P=$3 - in the long run

QS=QD = 2,600,000 - 200,000x3 = 2,000,000 (bushels)

Farms = 2,000,000 (bushels) / 1000 (bushels/farm) = 2000